How many kilograms of oxygen are required to form 2.86kg of CU2O?
1
Expert's answer
2020-08-11T04:35:30-0400

4Cu + O2 → 2Cu2O

M(O2) = 32 g/mol

M(Cu2O) = 143 g/mol

n=m/M

n(Cu2O) = 2860 g / (143 g/mol) = 20 mol

Proportion:

1 mol – 2 mol

x mol – 20 mol

x = 10 mol (O2)

m=n×M

m(O2) = 10 mol × 32 g/mol = 320 g

Answer: 0.32 kg